# thanks: corresponding author

### v2

@@ V.3 Interface between the annulus and the Morse spacetime
$\mathrm{d}s^{2}_{\mathrm{N}}=\mathrm{d}\sigma^{2}+\alpha^{2}(\sigma)\big{[}\mathrm{d}\psi^{2}+\sin^{2}(\psi)\,\mathrm{d}\Omega_{2}^{2}\big{]}$ (36)
$+\beta^{2}(\sigma)\left(\cos(2\psi)\,\mathrm{d}\sigma-\sigma\sin(2\psi)\,\mathrm{d}\psi\right)^{2},$ (37)
@@
$a(\sigma_{1})=\xi_{1},\qquad b(\sigma_{1})=1.$ (39)
@@
$\dot{a}(\sigma_{1})=1,\qquad\dot{b}(\sigma_{1})=0.$ (40)
@@ V.4 Homotopy between the vector fields along the neck
This assumption simplifies our analysis by eliminating potential complications in the deformation of the vector fields; in particular, because the underlying Riemannian metric is fixed, the conditions (25)–(26) and $\bigl{[}g^{R}_{\mu\nu,\alpha}\bigr{]}=0$ are automatically satisfied. Consequently, the only remaining task is to continuously deform the line field $\bigl{.}k^{-}_{\mu}\bigr{|}_{S^{3}}$—with its radial component reversed, since the gluing diffeomorphism is orientation-reversing—into the Morse line field $\bigl{.}v^{-}_{\mu}\bigr{|}_{S^{3}}$ while keeping both the fields and their derivatives smooth.
@@
with $c:[\eta_{-1/2},\eta_{0}]\to[0,1]$, $c(\eta_{-1/2})=1$, $c(\eta_{0})=0$. This deforms $w_{1,r}^{\mu}$ to $s^{\mu}$. Decompose $s^{\mu}$ further as
@@ Acknowledgements.
A. P., S. A., L. H. K., and S. L. thank Alexandros Kehagias for valuable discussions. A. P. also thank José M.M. Senovilla, Tina Harriott, and Jeff G. Williams for their insightful suggestions. B. S. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2024-04063.

### v3

@@ V.3 Interface between the annulus and the Morse spacetime
$\mathrm{d}s^{2}_{\mathrm{N}}=\mathrm{d}\sigma^{2}+\alpha^{2}(\sigma)\big[\mathrm{d}\psi^{2}+\sin^{2}(\psi)\,\mathrm{d}\Omega_{2}^{2}\big]$
$+\beta^{2}(\sigma)\left(\cos(2\psi)\,\mathrm{d}\sigma-\sigma\sin(2\psi)\,\mathrm{d}\psi\right)^{2},$
@@
$\alpha(\sigma_{1})=\xi_{1},\qquad\beta(\sigma_{1})=1.$ (37)
@@
$\dot{\alpha}(\sigma_{1})=1,\qquad\dot{\beta}(\sigma_{1})=0.$ (38)
@@ V.4 Homotopy between the vector fields along the neck
This assumption simplifies our analysis by eliminating potential complications in the deformation of the vector fields; in particular, because the underlying Riemannian metric is fixed, the conditions (25) and (26) and $\bigl[g^{R}_{\mu\nu,\alpha}\bigr]=0$ are automatically satisfied. Consequently, the only remaining task is to continuously deform the line field $\bigl.k^{-}_{\mu}\bigr|_{S^{3}}$—with its radial component reversed, since the gluing diffeomorphism is orientation-reversing—into the Morse line field $\bigl.v^{-}_{\mu}\bigr|_{S^{3}}$ while keeping both the fields and their derivatives smooth.
@@
with $c:[\eta_{-1/2},\eta_{0}]\to[0,1]$, $c(\eta_{-1/2})=1$, and $c(\eta_{0})=0$. This deforms $w_{1,r}^{\mu}$ to $s^{\mu}$. Decompose $s^{\mu}$ further as
@@ Acknowledgements.
A. P., S. A., L. H. K., and S. L. thank Alexandros Kehagias for valuable discussions. A. P. also thanks José M.M. Senovilla, Tina Harriott, and Jeff G. Williams for their insightful suggestions. B. S. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2024-04063.